OpenStudy (anonymous):

What is the solution set of |x – 6| = 10? 16 or –4 16 or –16 16 16 or 4

OpenStudy (anonymous):

i think its a

OpenStudy (anonymous):

or d

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a

OpenStudy (anonymous):

So its a!

Parth (parthkohli):

Absolute value of x - 6 = 10 means that x - 6 is 10 units away from 0.

OpenStudy (anonymous):

just put in eq you will see

OpenStudy (anonymous):

Ok! Thanks!

Parth (parthkohli):

\( \color{Black}{\Rightarrow x - 6 = \pm 10 }\) \( \color{Black}{\Rightarrow x = 16, -4 }\)

OpenStudy (anonymous):